# Slack-Pack algorithm for Meir-Moser packing problem

### v1

@@ Abstract
Keywords: square packing, Meir–Moser problem, harmonic series.
Mathematics Subject Classification: 52C15, 05B40.
@@ 1 Introduction
Note that problems on quadrating a rectangle are well known in combinatorics [1]. One of their traditional interpretations consists in constructing an electrical circuit, which transmits the unit electric current between two terminals with the potential difference of $\pi^{2}/6-1$ with the help of infinitely many wires, whose conductivities equal $1/2,1/3,1/4,\ldots$. However, in this case, as distinct from the Tutte problem, the way in which we can apply Kirchhoff laws is not clear. As a base we use an algorithm, which, in our opinion, solves the packing problem for infinitely many details.
There are several algorithms for solving the Meir-Moser problem. M.M. Paulhus [10] has applied one variant of the greedy algorithm, which implies that for packing a current detail, one chooses an empty rectangular box with the minimal admissible width (if the number of such boxes is more than one, then one chooses the box with the minimal height). Then one places the detail in a corner of the chosen box and performs the first cut. This cut (performed along the detail edge parallel to the lesser box side) results in dividing the box onto two parts (see the second illustration in Fig. 1). As a result of the second cut performed in one of these parts, one finally gets the cut-out detail and two new boxes. These boxes can be further used for packing details [^1] Clearly, if the detail and the initial box have the same sizes, then no new boxes appear..
Fig. 1: Illustration of the work of the Paulhus algorithm in packing details in a chosen box
@@
Fig. 2: Illustration of the work of the Paulhus algorithm in packing first 15 rectangles with side lengths of $\frac{1}{n}\times\frac{1}{n+1}$, $n\in\mathbb{N}$
@@
Fig. 3: Illustration of the work of the Wästlund algorithm in packing first 195 $\frac{1}{n}\times\frac{1}{n+1}$ rectangles, $n\in\mathbb{N}$, $n_{0}=100$
Note that authors of the paper [2] and subsequent papers [7][4], and [5] studied the perfect packing problem for the square sheet, whose area equals $\zeta(2t)$, and square details, whose side length equals $1/n^{t}$, where $n$ ranges from 1 to infinity. In particular, the existence of such packing was proved by A. Chalcraft [2] for $t\in[0.5864,0.6]$ and by A. Joós [7] for $t\in[\log_{3}2,2/3]$, $\ \log_{3}2\approx 0.63$; J. Januszewski and L. Zielonka [4] have extended this result for $t\in(1/2,2/3)$ and thus obtained a result stronger than that described in [12]. Finally, recently Janusz Januszewski [5] has proposed a very simple specific algorithm for proving the existence of a perfect packing for $t\in(1/2,17/32]$.
@@
The further part of the paper has the following structure. In Section 2, we introduce the main terms, describe a packing algorithm, and state the main results. In Section 3, we prove lemmas on the unification of sizes of boxes obtained in the stacking process and on their relative area. Section 4 is devoted to lemmas on the relative area of the rest boxes and on the final time moment of the first stage of the algorithm. In Section 5, we prove the main result of this paper. In Conclusion, we describe results of numerical experiments. In Appendix, we prove one auxiliary result on uniformly distributed random values; we need it for justifying our assumption.
@@ 2.1 Main terms and denotations
Fig. 4: Illustration of the packing of a row of details in a stripe by the stacking method; obtained normal boxes are colored in yellow and endpoints are colored in green
@@ 2.2 The Slack-Pack algorithm
Fig. 5: Illustration of the work of the Slack-Pack algorithm in packing first 35 details from the set $S_{n}$, $n_{0}=34$, $\gamma=10/7$. Endpoints and normal boxes are colored in green and yellow, correspondingly
@@
Fig. 6: Illustration of the work of the Slack-Pack algorithm in packing first 83 details from the set $S_{n}$, $n_{0}=34$, $\gamma=10/7$. Endpoints and normal boxes are colored in shades of green and yellow, correspondingly
@@ 2.3 The main theorem
Assume that the width of the endpoint obtained at time moment $n$ represents a random value $\xi_{n}$ uniformly distributed on the segment, whose left and right boundaries have, correspondingly, the order of the minimal and maximal possible width of an endpoint at time moment $n$, while the order of the height of this endpoint is $1/n$. Note that at time moment $t$ we have a set of endpoints obtained at time moments $n$, $n<t$. Assume that they obey a probability distribution law which agrees with all previously made assumptions. In Appendix, we prove that for such a mixture of uniformly distributed random values $\xi_{n}$, the asymptotics of the mean value of the expression $(1/n)/\xi_{n}$ is $(\gamma^{2}-1)\ln(t)$. We also adduce results of numerical experiments which with various $\gamma$ prove that the following ratio of the sample average to $\ln(t)$ tends to $\gamma^{2}-1$:
$\frac{1}{|{\mathcal{B}}_{ep,2}(t)|\ln(t)}\sum_{B\in{\mathcal{B}}_{ep,2}(t)}\frac{h(B)}{w(B)}.$
We also demonstrate that values of the mixture parameters used by us agree with experimental data.
Therefore, in our opinion, the following assumption is quite natural.
@@
$\frac{1}{|{\mathcal{B}}_{ep,2}(t)|}\sum_{B\in{\mathcal{B}}_{ep,2}(t)}\frac{h(B)}{w(B)}<O(\ln(t)).$ (6)
@@
Let assumptions 1 and 2 be fulfilled, and $\gamma\in(\sqrt{3/2},3/2)$. Then for any $\delta>0$, with sufficiently large $n_{0}$ for all $t\in T_{crit}$ inequality (2) is fulfilled and, consequently, it is possible to obtain a perfect packing by using the Slack-Pack algorithm.
@@ 3 Lemmas on normal boxes
According to results of a more accurate analysis, the maximal value of $h(B_{n})^{\gamma}/w(B_{n})$ for $B_{n}\in{\mathcal{B}}_{norm,1}$ is attained at the end of the first row in the normal box.
@@ 4.1 Lemmas on the ratio of endpoints of the first kind
Recall that endpoints of the first kind can be obtained by packing a row of details in the active box, which represents a stripe or some other endpoint of the first kind. Denote by ${\mathcal{B}}_{s-ep,1}(t)$ the set of all endpoints of the first kind obtained in a stripe up to time moment $t$. Then ${\mathcal{B}}^{\prime}_{s-ep,1}(t)=\bigcup\limits_{n=n_{0}}^{t}{\mathcal{B}}_{s-ep,1}(n)$.
@@
If the active box represents an endpoint of the first kind, then by packing a row of details in it we get a new endpoint of the first kind. Evidently, its height does not exceed the height of the initial endpoint. By Lemma 5 we conclude that
$\sum_{B\in{\mathcal{B}}_{ep,1}(t)}h(B)\leq\sum_{B\in{\mathcal{B}}^{\prime}_{s-ep,1}(t)}h(B)<p,$
@@ 4.2 Lemmas on the starting time moment of the second stage of the algorithm
Choosing an endpoint of the first kind as the active box and then packing a row of details in it, we get a new endpoint of the first kind and normal boxes of the second kind.222for example, in Fig. 6, by packing the row of details $D_{113}-D_{114}$ in the endpoint above detail $D_{50}$ we get a new endpoint of the first kind and 2 normal boxes of the second kind Evidently, the total area of new boxes is less than the area of the initial endpoint. Since with $t\leq t_{0}$ normal boxes of the second kind can be obtained only by packing details in endpoints of the first kind, we conclude that
$S_{norm,2}(t)+S_{ep,1}(t)\leq\sum_{B\in{\mathcal{B}}^{\prime}_{s-ep,1}(t)}S(B).$
@@
$\sum_{B\in{\mathcal{B}}^{\prime}_{s-ep,1}(t)}h(B)<p=2\sqrt{S_{com}(n_{0})}=O(n_{0}^{-1/2}).$
@@
$S_{norm,2}(t)+S_{ep,1}(t)=\sum_{B\in{\mathcal{B}}^{\prime}_{s-ep,1}(t)}h(B)w(B)=O(n_{0}^{-3/2}).$
@@
Consider the first stripe cut out from the LRP.333in Fig. 6, this is the stripe, where the row of details $D_{34}-D_{39}$ is packed Let the symbol $k_{0}$ stand for the quantity of details in this stripe and let ${\widetilde{B}}$ do for the normal box obtained by packing the last detail in this row.444in Fig. 6, this is the normal box located above detail $D_{39}$ Then the index of this normal box is $n_{0}+k_{0}-1$. Evidently, the second stage will start not later the time moment, when the row of details is packed in ${\widetilde{B}}$.
@@ 6 Conclusion
We have also performed certain experiments for verifying inequality (4) in Assumption 1. For the value $1+\sigma$ we used the expression $h({\widetilde{B}})^{\gamma}/w({\widetilde{B}})$, where ${\widetilde{B}}$ denotes the last normal box in the first stripe. Experiments were performed for sets of details $R_{n}$ with the same values of $n_{0}$ and $\gamma$ as those used in previous experiments. According to obtained results, inequality (4) is fulfilled for all considered normal boxes.
See the Appendix for theoretical and experimental arguments in favor of Assumption 2. See https://github.com/AlexeyKislovskiy/Slack-Pack-algorithm for the code of the program that visualizes the work of various packing algorithms and https://github.com/sensen65/Slack-Pack-algorithm-Meir-Moser for the code of the program that implements the Slack-Pack algorithm for $10^{10}$ details.
@@
Appendix. About Assumption 2
@@
The probability distribution function $F(x)$ of the random value $\xi_{n}$, which is uniformly distributed on the segment $[1/n^{\gamma},1/t]$, takes (on this segment) the following form:
@@
Acknowledgements
The authors is grateful to Olga Kashina for her help in the translation of this paper.

### v2

@@ Abstract
Mathematics Subject Classification: 52C15, 05B40
Keywords: Rectangle packing, Square packing, Meir–Moser problem, Harmonic series.
@@ 1 Introduction
Note that problems on quadrating a rectangle are well known in combinatorics [1]. One of their traditional interpretations consists in constructing an electrical circuit, which transmits the unit electric current between two terminals with the potential difference of $\pi^{2}/6-1$ with the help of infinitely many wires, whose conductivities equal $1/2,1/3,1/4,\ldots$. However, in this case, as distinct from the Tutte problem, the way in which we can apply Kirchhoff laws is not clear.
There are several algorithms for solving the Meir-Moser problem. M. Paulhus (see [10]) has applied one variant of the greedy algorithm, which implies that for packing a current detail, one chooses an empty rectangular box with the minimal admissible width (if the number of such boxes is more than one, then one chooses the box with the minimal height). Then one places the detail in a corner of the chosen box and performs the first cut. This cut (performed along the detail edge parallel to the lesser box side) results in dividing the box onto two parts (see the second illustration in Fig. 1). As a result of the second cut performed in one of these parts, one finally gets the cut-out detail and two new boxes. These boxes can be further used for packing details [^1] Clearly, if the detail and the initial box have the same sizes, then no new boxes appear..
Fig. 1: Illustration of the work of the Paulhus algorithm in packing details in a chosen box.
@@
Fig. 2: Illustration of the work of the Paulhus algorithm in packing first 15 rectangles with side lengths of $\frac{1}{n}\times\frac{1}{n+1}$, $n\in\mathbb{N}$.
@@
Fig. 3: Illustration of the work of the Wästlund algorithm in packing first 195 $\frac{1}{n}\times\frac{1}{n+1}$ rectangles, $n\in\mathbb{N}$, $n_{0}=100$.
Note that authors of the paper [2] and subsequent papers [7][5], and [4] studied the perfect packing problem for the square sheet, whose area equals $\zeta(2t)$, and square details, whose side length equals $1/n^{t}$, where $n$ ranges from 1 to infinity. In particular, the existence of such packing was proved by A. Chalcraft [2] for $t\in[0.5864,0.6]$ and by A. Joós [7] for $t\in[\log_{3}2,2/3]$, $\ \log_{3}2\approx 0.63$; J. Januszewski and L. Zielonka [5] have extended this result for $t\in(1/2,2/3)$ and thus obtained a result stronger than that described in [12]. Finally, recently J. Januszewski [4] has proposed a very simple specific algorithm for proving the existence of a perfect packing for $t\in(1/2,17/32]$.
@@
The further part of the paper has the following structure. In Section 2, we introduce the main terms, describe a packing algorithm, and state the main results. In Section 3, we prove lemmas on the unification of sizes of boxes obtained in the stacking process and on their relative area. Section 4 is devoted to lemmas on the relative area of the rest boxes and on the final time moment of the first stage of the algorithm. In Section 5, we prove the main result of this paper. In Conclusion, we describe results of numerical experiments. In Appendix A, we prove one auxiliary result on uniformly distributed random values; we need it for justifying our assumption.
@@ 2.1 Main terms and denotations
Fig. 4: Illustration of the packing of a row of details in a stripe by the stacking method; obtained normal boxes are colored in yellow and endpoints are colored in green.
@@ 2.2 The Slack-Pack algorithm
Fig. 5: Illustration of the work of the Slack-Pack algorithm in packing first 35 details from the set $S_{n}$, $n_{0}=34$, $\gamma=10/7$. Endpoints and normal boxes are colored in green and yellow, correspondingly.
@@
Fig. 6: Illustration of the work of the Slack-Pack algorithm in packing first 83 details from the set $S_{n}$, $n_{0}=34$, $\gamma=10/7$. Endpoints and normal boxes are colored in shades of green and yellow, correspondingly.
@@ 2.3 The main theorem
Assume that the width of the endpoint obtained at time moment $n$ represents a random value $\xi_{n}$ uniformly distributed on the segment, whose left and right boundaries have, correspondingly, the order of the minimal and maximal possible width of an endpoint at time moment $n$, while the order of the height of this endpoint is $1/n$. Note that at time moment $t$ we have a set of endpoints obtained at time moments $n$, $n<t$. Assume that they obey a probability distribution law which agrees with all previously made assumptions. In Appendix A, we prove that for such a mixture of uniformly distributed random values $\xi_{n}$, the asymptotics of the mean value of the expression $(1/n)/\xi_{n}$ is $(\gamma^{2}-1)\ln(t)$. We also adduce results of numerical experiments which with various $\gamma$ prove that the following ratio of the sample average to $\ln(t)$
$\frac{1}{\ln(t)\,|{\mathcal{B}}_{ep,2}(t)|}\sum_{B\in{\mathcal{B}}_{ep,2}(t)}\frac{h(B)}{w(B)}.$
tends to $\gamma^{2}-1$. We also demonstrate that values of the mixture parameters used by us agree with experimental data.
Therefore, the following assumption is quite natural.
@@
$\frac{1}{|{\mathcal{B}}_{ep,2}(t)|}\sum_{B\in{\mathcal{B}}_{ep,2}(t)}\frac{h(B)}{w(B)}=O(\ln(t)).$ (6)
@@
Let assumptions 1 and 2 be fulfilled for some fixed $\gamma$, $\gamma\in(\sqrt{3/2},3/2)$. Then (for this $\gamma$) for any $\delta>0$, with sufficiently large $n_{0}$ for all $t\in T_{crit}$ inequality (2) is fulfilled and, consequently, it is possible to obtain a perfect packing by using the Slack-Pack algorithm.
@@ 3 Lemmas on normal boxes
According to results of a more accurate analysis, the maximal value of $\frac{h(B_{n})^{\gamma}}{w(B_{n})}$ for $B_{n}\in{\mathcal{B}}_{norm,1}$ is attained at the end of the first row in the normal box.
@@ 4.1 Lemmas on the ratio of endpoints of the first kind
Recall that endpoints of the first kind can be obtained by packing a row of details in the active box, which represents a stripe or some other endpoint of the first kind. Denote by ${\mathcal{B}}_{s.ep,1}(t)$ the set of all endpoints of the first kind obtained in a stripe up to time moment $t$. Then ${\mathcal{B}}^{\prime}_{s.ep,1}(t)=\bigcup\limits_{n=n_{0}}^{t}{\mathcal{B}}_{s.ep,1}(n)$.
@@
If the active box represents an endpoint of the first kind, then by packing a row of details in it we get a new endpoint of the first kind. Evidently, its height does not exceed the height of the initial endpoint (see  Fig. 7). By Lemma 6 we conclude that
$\sum_{B\in{\mathcal{B}}_{ep,1}(t)}h(B)\leq\sum_{B\in{\mathcal{B}}^{\prime}_{s.ep,1}(t)}h(B)<p,$
@@ 4.2 Lemmas on the starting time moment of the second stage of the algorithm
Choosing an endpoint of the first kind as the active box and then packing a row of details in it, we get a new endpoint of the first kind and normal boxes of the second kind. [^2] For example, in Fig. 6, by packing the row of details $D_{113}-D_{114}$ in the endpoint above detail $D_{50}$ we get a new endpoint of the first kind and 2 normal boxes of the second kind. Evidently, the total area of new boxes is less than the area of the initial endpoint. Since with $t\leq t_{0}$ normal boxes of the second kind can be obtained only by packing details in endpoints of the first kind, we conclude that
$S_{norm,2}(t)+S_{ep,1}(t)\leq\sum_{B\in{\mathcal{B}}^{\prime}_{s.ep,1}(t)}S(B).$
@@
$\sum_{B\in{\mathcal{B}}^{\prime}_{s.ep,1}(t)}h(B)<p=2\sqrt{S_{com}(n_{0})}=O(n_{0}^{-1/2}).$
@@
$S_{norm,2}(t)+S_{ep,1}(t)=\sum_{B\in{\mathcal{B}}^{\prime}_{s.ep,1}(t)}h(B)w(B)=O(n_{0}^{-3/2}).$
@@
Consider the first stripe cut out from the LRP. [^3] In Fig. 6, this is the stripe, where the row of details $D_{34}-D_{39}$ is packed. Let the symbol $k_{0}$ stand for the quantity of details in this stripe and let ${\widetilde{B}}$ do for the normal box obtained by packing the last detail in this row. [^4] In Fig. 6, this is the normal box located above detail $D_{39}$. Then the index of this normal box is $n_{0}+k_{0}-1$. Evidently, the second stage will start not later the time moment, when the row of details is packed in ${\widetilde{B}}$.
@@ 6 Conclusion
We have also performed certain experiments for verifying inequality (4) in Assumption 1. For the value $1+\sigma$ we used the expression $h({\widetilde{B}})^{\gamma}/w({\widetilde{B}})$, where ${\widetilde{B}}$ denotes the last normal box in the first stripe. Experiments were performed for sets of details $R_{n}$ with the same values of $n_{0}$ and $\gamma$ as those used in previous experiments. According to obtained results, inequality (4) is fulfilled for all considered normal boxes. See also the Appendix A for theoretical and experimental arguments in favor of Assumption 2.
@@
Acknowledgements. The authors is grateful to Olga Kashina for her help in the translation of this paper.
Data Availability. The data is publicly available in public repositories. See https://github.com/AlexeyKislovskiy/Slack-Pack-algorithm for the code of the program that visualizes the work of various packing algorithms. See also the code of the program in https://github.com/sensen65/Slack-Pack-algorithm-Meir-Moser that implements the Slack-Pack algorithm for $10^{10}$ details.
Declarations
Conflict of interest. The authors have no conflicts of interest relevant to the content of this article.
Appendix A About Assumption 2
@@
The cummulative density function $F(x)$ of the random value $\xi_{n}$, which is uniformly distributed on the segment $[1/n^{\gamma},1/t]$, takes (on this segment) the following form: